OpenStudy (anonymous):

what is the oxidation number of Mg in MgSO4

OpenStudy (anonymous):

\[Mg ^{+2} +SO _{4}^{-2} \rightarrow MgSO _{4}\] \[Mg + (1\times biloks \S) + (4 \times biloks O) = 0\] since S is free element so the oxidation number is zero number oxidation for O is -2 so \[Mg+(1 \times 0)+ (4 \times (-2) =0\] \[number oxidation of Mg becomes Mg = +8\]